OpenStudy (anonymous):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin r = 3 - 4 sin θ

OpenStudy (mathmale):

Tricky question. You might learn a lot from actually graphing r=3-4sin theta. Write r as \[r(\theta)=3-4\sin \theta\]

OpenStudy (mathmale):

How does the fact that that '3' in this formula is always the same, whereas -4sin theta could be positive or negative, depending upon the value of the angle theta. If \[r(-\theta)=r(\theta), \]

OpenStudy (mathmale):

the function is EVEN. Hint: is sin x an even function or an odd one?

OpenStudy (anonymous):

its not coming up when i graph it

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

If \[r(-\theta)=-r(\theta), \]

OpenStudy (mathmale):

this function is ODD.

OpenStudy (mathmale):

If neither equation is true, then the function neither even nor odd.

OpenStudy (anonymous):

Why won't it come up on my calculator graph @mathmale

OpenStudy (mathmale):

that's a separate question, Robert. Do you want to discuss the question you've posted or do you want to discuss use of your calculator?

OpenStudy (anonymous):

@mathmale well would it be easier to just graph it then see where it is symmetric

OpenStudy (mathmale):

sure, but that won't help you if your teacher expects y ou to determine symmetry (or the lack of it) without a calculator.

OpenStudy (mathmale):

What kind of calc do you own?

OpenStudy (anonymous):

TI 84 plus @mathmale